OpenStudy (anonymous):

Mehrab works in a department store selling sports equipment. Baseball gloves cost $29 each and bats cost $14 each. In one shift, he sells 28 items. His receipts total is $647. How many bats did he sell? How many gloves? I understand a bit of the question.

OpenStudy (anonymous):

Let X be gloves and let y be bats. I get that. and i came up with my first equation, 29x+14y=647

OpenStudy (anonymous):

you could start with \[x+y=28\\ 29x+14y=647\] if you like

OpenStudy (anonymous):

that looks good

OpenStudy (anonymous):

oh, OK now I get it. Thank you, @satellite73. I just could not come up with a second equation. now I know what to do!